# StyleDubber: Towards Multi-Scale Style Learning for Movie Dubbing

### v2

@@ 4.3 Evaluation Metrics
Objective metrics. To measure whether the generated speech carries the desired speaker identity and emotion, speaker identity similarity (SPK-SIM) and emotion accuracy (EMO-ACC) are employed as in Chen et al. (2022). We adopt the Mel Cepstral Distortion Dynamic Time Warping (MCD-DTW) to measure the difference between generated speech and real speech. We also adopt the metric MCD-DTW-SL, which is MCD-DTW weighted by duration consistency Chen et al. (2022). Besides, the Word Error Rate (WER) is used to measure pronunciation accuracy by the publicly available whisper Radford et al. (2023) as the ASR model.
Subjective metrics. We also provide subjective evaluation results via conducting a human study using a 5-scale mean opinion score (MOS) in two aspects: naturalness and similarity. Following the settings in Chen et al. (2022), all participants are asked to assess the sound quality of 25 randomly selected audio samples from each test set.
@@ 4.4 Performance Evaluations
Dub 1.0 (Original setting Chen et al. (2022)) Ground-truth speaker in test set 2,779
@@
Dub 3.0 (Unseen speaker setting) Unseen speaker 33$\times$153
@@
We evaluate our method in three experimental settings as shown in Table 3. The first setting is the same as in Chen et al. (2022), which uses ground-truth audio as reference audio. However, this is impractical in real-world applications. Thus, we design two new and more reasonable settings: “Dub 2.0” uses non-ground truth audio of the same speaker as reference audio; “Dub 3.0” uses the audio of unseen characters (from another dataset) as reference audio. We compare with six recent related baselines to comprehensively analyze. More details on baselines are in Appendix B. Furthermore, we will release the detailed configuration for all experiment settings for the GRID and V2C Animation datasets.
@@
Results under Dub 2.0 setting. We report the V2C results in Table 2 and put the results on GRID in Appendix due to page limit. Despite Dub 2.0 is much more challenging than 1.0, our method still outperforms other methods on all metrics. The SPK-SIM and WER significantly improve 48.29% and 9.15% than previous best results. In terms of EMO-ACC, our method achieves relative 3.38% improvements, reflecting the ability to maintain emotions with video, although providing a reference voice that may be inconsistent with the emotion of the current scene. Additionally, the MCD-DTW is improved by 6.11%, which indicates our method achieves minimal acoustic difference even in challenging setting 2.0. Furthermore, the lowest MCD-DTW-SL shows that our method achieves almost the same duration sync as the groud-truth video. Finally, the human subjective evaluation results (see MOS-N and MOS-S) also show that our StyleDubber can generate speeches that are closer to realistic speech in both naturalness and similarity.
@@
Fastspeech2 Ren et al. (2021) X 2.91 $\pm$ 0.13 3.02 $\pm$ 0.09 21.11 27.73
StyleSpeech Min et al. (2021) X 3.17 $\pm$ 0.06 3.22 $\pm$ 0.15 55.81 0.9340
Zero-shot TTS Zhou et al. (2022) X 3.53 $\pm$ 0.12 3.35 $\pm$ 0.07 57.23 31.47
Fastspeech2* Ren et al. (2021) 2.97 $\pm$ 0.12 3.03 $\pm$ 0.29 26.79 30.27
StyleSpeech* Min et al. (2021) 3.31 $\pm$ 0.18 3.22 $\pm$ 0.10 58.71 105.64
Zero-shot TTS* Zhou et al. (2022) 3.62 $\pm$ 0.09 3.31 $\pm$ 0.13 61.12 35.10
V2C-Net Chen et al. (2022) 3.05 $\pm$ 0.07 2.83 $\pm$ 0.09 38.43 143.54
HPMDubbing Cong et al. (2023) 3.11 $\pm$ 0.08 2.92 $\pm$ 0.09 49.31 106.45
Face-TTS Lee et al. (2023) 3.10 $\pm$ 0.05 3.17 $\pm$ 0.15 33.80 231.63
Ours 3.94$\pm$ 0.12 3.87$\pm$ 0.14 71.85 25.91
@@ 4.5 Qualitative Results
We visualize the mel-spectrogram of reference audio, ground-truth audio, and synthesized audios by ours and the other two state-of-the-art methods in Figure 3. We highlight the regions in mel-spectrograms where significant differences are observed among these methods in reconstruction details (green boxes), and pause (blue boxes), respectively. We find that our method is more similar to the ground-truth mel-spectrogram, which has clearer and distinct horizontal lines in the spectrum to benefit the fine-grained pronunciation expression of speakers (see green box). By observing the blue boxes, we find that our method can learn nature pauses to achieve better sync by aligning phonemes and lip motion.
@@
1 w/o MPS 60.94 77.26 9.90 9.98 41.18
2 w/o USL 51.97 47.07 10.23 10.43 42.77
3 w/o PLA 58.92 80.77 9.59 11.47 45.48
4 Quasi-phoneme $v.s.$ frame 54.26 80.76 9.75 9.84 43.83
5 w/o $R_{A\rightarrow L}$ 60.27 77.79 9.81 9.90 42.61
6 w/o $Z_{V\rightarrow L}$ 55.05 79.30 9.82 9.91 41.49
7 w/o U-MelDecoder 51.75 47.31 10.31 10.41 42.80
8 w/o U-Post 51.58 80.79 9.52 9.61 44.33
9 Full model 49.83 82.24 9.44 9.53 45.51
Table 5: Ablation study of the proposed method on the V2C benchmark dataset with 2.0 setting, respectively.
@@ 4.6 Ablation Studies
To further study the influence of the individual components in StyleDubber, we perform the comprehensive ablation analysis using the version 2.0 of V2C.
@@
Effectiveness of U-MelDecoder and U-post. To prove the effect of each module in USL, we remove the utterance-level style learning on mel-decoder and post-net, respectively. As shown in Row 8-9 of Table 5, when removing the U-post, the performance also drops but is not as large as removing the U-MelDecoder. This indicates that U-MelDecoder is critical to the generation of spectrum, while U-post only works on refining spectrum in 80 channels so that the impact is relatively small.
@@ 7 Ethics Statement
  • In Section A, we analyze the challenging of the V2C benchmark compared with the traditional TTS benchmark.
@@
  • In Section C, we report the Grid result on Dub setting 2.0.
@@ Appendix A The challenges in V2C benchmark
The V2C benchmark significantly differs from traditional TTS benchmark , and it is more challenging in the following aspects: (1) The data scale of V2C dataset is much smaller in terms of either the number of data items or speech length (see Figure 4 (a)-(b)). There are only 9374 video clips in V2C, and most of its audio is shorter than 5s. In contrast, FS2 and Stylespeech are trained on LJspeech and LibriTTS with 13,100 and 149,753 samples, most of which are longer than 5s. Although LJspeech also looks similar in size to V2C, it is a single-speaker dataset, so V2C allocates fewer samples to each speaker. (2) V2C has the largest variance of pitch compared to TTS tasks due to exaggerated expressions of cartoon characters (see Figure 1 (c) and more details in Tab. 2 of V2C-Net). (3) The audio of V2C contains background noise or music, like car whistle and alarm clock sound, et al. Signal-to-noise Ratio (SNR) of V2C is the lowest (Figure 4 (d)). In summary, unlike the large-scale clean TTS datasets, V2C is much more challenging, and the well-known TTS methods suffer performance degradation.
@@ Appendix B Baselines
We compare our method against six closely related methods with available codes. 1) StyleSpeech Min et al. (2021) is a multi-speaker voice clone method that synthesizes speech in the style of the target speaker via meta-learning; 2) FastSpeech2 Ren et al. (2021) is a popular multi-speaker TTS method for explicitly modeling energy and pitch in speech; 3) Zero-shot TTS Zhou et al. (2022) is a content-dependent fine-grained speaker method for zero-shot speaker adaptation. 4) V2C-Net Chen et al. (2022) is the first visual voice cloning model for movie dubbing; 5) HPMDubbing Cong et al. (2023) is a hierarchical prosody modeling for movie dubbing, which bridges video representations and speech attributes from three levels: lip, facial expression, and scene. 6) Face-TTS Lee et al. (2023) is a novel face-styled speech synthesis within a diffusion model, which leverages face images to provide a robust characteristic of speakers. In addition, for a fair comparison, for the pure TTS method, we adopt the setting as Chen et al. (2022), which takes video embedding as an additional input, before the duration predictor to predict the duration.
GT - 100.00 21.41 - 0.0 0.0 4.85 $\pm$ 0.05 4.90 $\pm$ 0.03
GT Mel + Vocoder - 97.57 22.41 - 4.11 4.16 4.72 $\pm$ 0.08 4.76$\pm$ 0.10
Fastspeech2 Ren et al. (2021) X 47.41 19.55 - 7.67 8.44 3.01$\pm$ 0.08 3.12$\pm$ 0.12
StyleSpeech Min et al. (2021) X 74.15 21.42 - 7.02 7.95 3.22 $\pm$ 0.06 3.31$\pm$ 0.016
Zero-shot TTS Zhou et al. (2022) X 82.25 19.35 - 6.21 6.76 4.11 $\pm$ 0.12 4.05 $\pm$ 0.09
StyleSpeech* Min et al. (2021) 74.52 23.35 - 7.01 7.82 3.38 $\pm$ 0.12 3.46$\pm$ 0.18
Fastspeech2* Ren et al. (2021) 59.58 19.82 - 7.24 7.96 3.31 $\pm$ 0.26 3.22 $\pm$ 0.10
Zero-shot TTS* Zhou et al. (2022) 81.34 21.05 - 6.27 7.29 4.18 $\pm$ 0.09 4.09 $\pm$ 0.14
V2C-Net Chen et al. (2022) 71.51 49.09 - 7.29 7.86 3.53 $\pm$ 0.07 3.39 $\pm$ 0.09
HPMDubbing Cong et al. (2023) 71.99 44.15 - 6.79 7.09 3.62 $\pm$ 0.10 3.47 $\pm$ 0.06
Face-TTS Lee et al. (2023) 34.14 39.15 - 7.77 8.59 3.42$\pm$ 0.18 3.39 $\pm$0.10
Ours 93.69 18.79 - 5.66 5.91 4.29 $\pm$ 0.11 4.12 $\pm$ 0.09
Table 6: The Grid results under Dub 2.0 setting, which uses non-ground truth of same character as reference audio. Since the Grid dataset does not have emotion labels, we did not calculate EMO-ACC.
Appendix C Grid result on Dub 2.0 setting
As discussed in Table 2, we also made the Dub 2.0 setting on the Grid dataset. Specifically, we use the same speaker from different video voices as reference audio to evaluate the generalization performance of the model. As shown in Table 6, we find that all models perform better on the Grid dataset, which may be due to the clean and simple speech environment. Although there have been improvements, our method still outperforms other methods on all metrics. Especially the speaker’s style, our models achieve the $93.69$% SPK-SIM, which reflects our StyleDubber ability to adapt to the speaker’s style through two-level learning. Additionally, the lowest WER proves the correctness of our pronunciation. Furthermore, the MCD-DTW is improved by 12.57% than the 2nd best method, which indicates our method achieves minimal acoustic difference even in challenging setting 2.0. Furthermore, the lowest MCD-DTW-SL shows that our method achieves better lip sync with the video content. In addition, the human subjective evaluation results, specifically the MOS-N (Mean Opinion Score-Naturalness) and MOS-S (Mean Opinion Score-Similarity), further confirm that our StyleDubber effectively produces speech outputs that closely resemble natural human speech in terms of both naturalness and similarity.

### v3

@@ 4.3 Evaluation Metrics
Objective metrics. To measure whether the generated speech carries the desired speaker identity and emotion, speaker identity similarity (SPK-SIM) is calculated by SECS Casanova et al. (2021), and emotion accuracy (EMO-ACC) is employed by pre-trained speech emotion recognition model Ye et al. (2023). Besides, we adopt the Mel Cepstral Distortion Dynamic Time Warping (MCD-DTW) to measure the difference between generated speech and real speech. We also adopt the metric MCD-DTW-SL, which is MCD-DTW weighted by duration consistency Chen et al. (2022a). The Word Error Rate (WER) Morris et al. (2004) is used to measure pronunciation accuracy by the publicly available whisper Radford et al. (2023) as the ASR model. Besides, we use the ASV model (WavLM-TDNN Chen et al. (2022b)) to comprehensively evaluate identity similarity (see Appendix D), following NaturalSpeech 3 Ju et al. (2024).
Subjective metrics. We also provide subjective evaluation results via conducting a human study using a 5-scale mean opinion score (MOS) in two aspects: naturalness and similarity. Following the settings in Chen et al. (2022a), all participants are asked to assess the sound quality of 25 randomly selected audio samples from each test set.
@@ 4.4 Performance Evaluations
Dub 1.0 (Original setting Chen et al. (2022a)) Ground-truth speaker in test set 2,779
@@
Dub 3.0 (Unseen speaker setting) Unseen speaker 4,851
@@
We evaluate our method in three experimental settings as shown in Table 3. The first setting is the same as in Chen et al. (2022a), which uses target audio as reference audio from test set. However, this is impractical in real-world applications. Thus, we design two new and more reasonable settings: “Dub 2.0” uses non-ground truth audio of the same speaker as reference audio; “Dub 3.0” uses the audio of unseen characters (from another dataset) as reference audio. We compare with six recent related baselines to comprehensively analyze. Furthermore, we will release the detailed configuration for all experiment settings for the GRID and V2C Animation datasets.
@@
Results under Dub 2.0 setting. We report the V2C results in Table 2. Despite Dub 2.0 is much more challenging than 1.0, our method still outperforms other methods on six metrics. The SPK-SIM and WER significantly improve 7.4% and 6.98% than the fastspeech 2 (TTS-textonly) and meta-stylespeech method, respectively. Additionally, the proposed method (StyleDubber) achieves the lowest MCD-DTW compared to all baselines, which indicates our method achieves minimal acoustic difference even in challenging setting 2.0. Furthermore, the lowest MCD-DTW-SL shows that our method achieves almost the same duration sync as the groud-truth video. Finally, the human subjective evaluation results (see MOS-N and MOS-S) also show that our StyleDubber can generate speeches that are closer to realistic speech in both naturalness and similarity.
@@
Fastspeech2 Ren et al. (2021) X 2.91 $\pm$ 0.13 3.02 $\pm$ 0.09 21.11 14.05
StyleSpeech Min et al. (2021) X 3.17 $\pm$ 0.06 3.22 $\pm$ 0.15 55.81 77.46
Zero-shot TTS Zhou et al. (2022) X 3.53 $\pm$ 0.12 3.35 $\pm$ 0.07 57.23 19.83
Fastspeech2* Ren et al. (2021) 2.97 $\pm$ 0.12 3.03 $\pm$ 0.29 26.79 18.38
StyleSpeech* Min et al. (2021) 3.31 $\pm$ 0.18 3.22 $\pm$ 0.10 58.71 89.11
Zero-shot TTS* Zhou et al. (2022) 3.62 $\pm$ 0.09 3.31 $\pm$ 0.13 61.12 19.25
V2C-Net Chen et al. (2022a) 3.05 $\pm$ 0.07 2.83 $\pm$ 0.09 38.43 112.71
HPMDubbing Cong et al. (2023) 3.11 $\pm$ 0.08 2.92 $\pm$ 0.09 49.31 106.81
Face-TTS Lee et al. (2023) 3.10 $\pm$ 0.05 3.17 $\pm$ 0.15 33.80 201.98
Ours 3.94$\pm$ 0.12 3.87$\pm$ 0.14 71.52 13.45
@@ 4.5 Qualitative Results
We visualize the mel-spectrogram of reference audio, ground-truth audio, and synthesized audios by ours and the other two state-of-the-art methods in Figure 3. We highlight the regions in mel-spectrograms where significant differences are observed among these methods in reconstruction details (green boxes), and pause (blue boxes), respectively. We find that our method is more similar to the ground-truth mel-spectrogram, which has clearer and distinct horizontal lines in the spectrum to benefit the fine-grained pronunciation expression of speakers (see green box). By observing the blue boxes, we find that our method can learn natural pauses to achieve better sync by aligning phonemes and lip motion.
@@
1 w/o MPA 39.74 77.26 9.90 9.98 41.18
2 w/o USL 32.81 47.07 10.23 10.43 42.77
3 w/o PLA 35.90 80.77 9.59 11.47 45.48
4 Quasi-phoneme $v.s.$ frame 33.62 80.76 9.75 9.84 43.83
5 w/o $R_{A\rightarrow L}$ 38.21 77.79 9.81 9.90 42.61
6 w/o $Z_{V\rightarrow L}$ 33.87 79.30 9.82 9.91 41.49
7 w/o U-MelDecoder 32.26 47.31 10.31 10.41 42.80
8 w/o U-Post 31.73 80.79 9.52 9.61 44.33
9 Full model 31.49 82.26 9.37 9.46 45.62
Table 5: Ablation study of the proposed method on the V2C benchmark dataset with 1.0 setting, respectively.
@@ 4.6 Ablation Studies
To further study the influence of the individual components in StyleDubber, we perform the comprehensive ablation analysis using V2C 1.0 version.
@@
Effectiveness of U-MelDecoder and U-post. To prove the effect of each module in USL, we remove the utterance-level style learning on mel-decoder and post-net, respectively. In other words, it still keeps an autoregressive manner by transformer-based decoder, and we just cut off the red arrow in Figure 2 (c). As shown in Row 7-8 of Table 5, when removing the U-post, the performance also drops but is not as large as removing the U-MelDecoder. This indicates that U-MelDecoder is critical to the generation of spectrum, while U-post only works on refining spectrum in 80 channels so that the impact is relatively small.
@@ 7 Ethics Statement
Acknowledgements
This work was supported in part by National Key R&D Program of China under Grant (2023YFB4502800), National Natural Science Foundation of China: 62322211, 61931008, 62236008, 62336008, U21B2038, 62225207, Fundamental Research Funds for the Central Universities (E2ET1104), “Pionee” and “Leading Goose” R&D Program of Zhejiang Province (2024C01023, 2023C01030). Yuankai Qi, Amin Beheshti, Anton van den Hengel, and Ming-Hsuan Yang are not sup- ported by the aforementioned fundings.
@@ Acknowledgements
  • In Section A, we analyze the challenges of the V2C benchmark compared with the traditional TTS benchmark.
@@
  • In Section C, we report the WER result by different Whisper versions on the V2C-Animation dataset.
  • In Section D, we report the speaker similarity result by the large WavLM-TDNN model on the V2C-Animation dataset and GRID dataset.
@@ Appendix A The challenges in V2C benchmark
The V2C benchmark significantly differs from traditional TTS benchmark , and it is more challenging in the following aspects: (1) The data scale of V2C dataset is much smaller in terms of either the number of data items or speech length (see Figure 4 (a)-(b)). There are only 9374 video clips in V2C, and most of its audio is shorter than 5s. In contrast, FS2 and Stylespeech are trained on LJspeech and LibriTTS with 13,100 and 149,753 samples, most of which are longer than 5s. Although LJspeech also looks similar in size to V2C, it is a single-speaker dataset, so V2C allocates fewer samples to each speaker. (2) V2C has the largest variance of pitch compared to TTS tasks due to exaggerated expressions of cartoon characters (see Figure 1 (c) and more details in Tab. 2 of V2C-Net). (3) The audio of V2C contains background noise or music, like car whistle and alarm clock sound, et al. Signal-to-noise Ratio (SNR) of V2C is the lowest (Figure 4 (d)). In summary, unlike the large-scale clean TTS datasets, V2C is much more challenging, and the well-known TTS methods suffer performance degradation. In this work, all experiments are conducted on the V2C denoise version. We will publish this version.
@@ Appendix B Baselines
We compare our method against six closely related methods with available codes. 1) StyleSpeech Min et al. (2021) is a multi-speaker voice clone method that synthesizes speech in the style of the target speaker via meta-learning; 2) FastSpeech2 Ren et al. (2021) is a popular multi-speaker TTS method for explicitly modeling energy and pitch in speech; 3) Zero-shot TTS Zhou et al. (2022) is a content-dependent fine-grained speaker method for zero-shot speaker adaptation. 4) V2C-Net Chen et al. (2022a) is the first visual voice cloning model for movie dubbing; 5) HPMDubbing Cong et al. (2023) is a hierarchical prosody modeling for movie dubbing, which bridges video representations and speech attributes from three levels: lip, facial expression, and scene. 6) Face-TTS Lee et al. (2023) is a novel face-styled speech synthesis within a diffusion model, which leverages face images to provide a robust characteristic of speakers. In addition, for a fair comparison, for the pure TTS method, we adopt the setting as Chen et al. (2022a), which takes video embedding as an additional input, before the duration predictor to predict the duration.
# Whisper’s Version WER on V2C-Animation (%) $\downarrow$
1 Whisper (base) 45.58
2 Whisper (large-v1) 23.88
3 Whisper (large-v2) 23.85
4 Whisper (large-v3) 22.55
Table 6: The WER test (the ground truth result) for various versions of Whisper on the V2C benchmark dataset.
Appendix C Whisper test on V2C-Animation dataset
In Table 6, the results show that large-v3 achieved the lowest WER, thus we re-selected large-v3 as the measurement tool to get more convincing results on the V2C-Animation dataset. Note that Whisper-Large V3 has not been fine-tuned in the V2C-Animation dataset, there is still some gap (WER in GT is 22.55 %), but it is enough to serve as a fair comparison. All results (Dub1.0, 2,0, and 3.0) still reflect that our StyleDubber is the best (see Table 1,2,4) and is more conducive to the clarity of movie dubbing. Considering the inference speed and computation memory, the Grid dataset still retains the original “Whisper-base” as the test benchmark. The “Whisper-base” achieves the 22.41 % GT WER on the GRID test as similar to the VDTTS Hassid et al. (2022) result in Table 2 (GRID evaluation).
Methods Visual Sim-O $\uparrow$ Sim-R $\uparrow$
Ground Truth - 0.79 n/a
Fastspeech2 Ren et al. (2021) X 0.10 0.19
StyleSpeech Min et al. (2021) X 0.14 0.23
Zero-shot TTS Zhou et al. (2022) X 0.12 0.21
Fastspeech2* Ren et al. (2021) 0.10 0.18
StyleSpeech* Min et al. (2021) 0.14 0.23
Zero-shot TTS* Zhou et al. (2022) 0.13 0.22
V2C-Net Chen et al. (2022a) 0.08 0.15
HPMDubbing Cong et al. (2023) 0.11 0.19
Face-TTS Lee et al. (2023) 0.09 0.12
Ours 0.25 0.34
Table 7: The V2C-Animation dataset results under the WavLM-TDNN similarity testing.
Appendix D WavLM-TDNN Similarity Testing
We employ the SOTA speaker verification model, WavLM-TDNN, to evaluate the speaker similarity between the prompt (i.e., the reference audio in the V2C task) and synthesized speech, following VALL-E Wang et al. (2023a), VoiceBox Le et al. (2023), and NaturalSpeech 3 Ju et al. (2024). WavLM-TDNN achieved the top rank at the VoxSRC Challenge 2021 and 2022 leaderboards and it is suitable as the SPK-SIM metric for the challenging V2C-nimation dataset Chen et al. (2022a). The similarity score predicted by WavLM-TDNN is in the range of [-1; 1], where a larger value indicates a higher similarity of input samples. Specifically, two metrics need to be calculated: (1) SIM-R represents the similarity with resynthesized audio, which is not comparable across models using different vocoders; (2) SIM-O is used to measure similarity against the original reference audio. Note that the report results based on the GE2E model (Table 1, 2, 4) are compared with the original waveform.
Methods Visual Sim-O $\uparrow$ Sim-R $\uparrow$
Ground Truth - 0.87 n/a
Fastspeech2 Ren et al. (2021) X 0.38 0.42
StyleSpeech Min et al. (2021) X 0.74 0.79
Zero-shot TTS Zhou et al. (2022) X 0.70 0.75
Fastspeech2* Ren et al. (2021) 0.48 0.52
StyleSpeech* Min et al. (2021) 0.74 0.79
Zero-shot TTS* Zhou et al. (2022) 0.69 0.74
V2C-Net Chen et al. (2022a) 0.43 0.55
HPMDubbing Cong et al. (2023) 0.46 0.56
Face-TTS Lee et al. (2023) 0.42 0.51
Ours 0.75 0.80
Table 8: The GRID dataset results under the WavLM-TDNN similarity testing.
As shown in Table 7,8, the results on two datasets show that even if the similarity measurement method is replaced, our StyleDubber still achieves the best performance in Sim-O and Sim-R. The result proves the effectiveness of StyleDubber, which proposes multi-scale style learning at phoneme and utterance levels and captures precise pronunciation in acoustic details and visual emotion for dubbing. Besides, we have several findings: (1) Compared with the GT result on the GRID dataset (0.87), the Sim-O result of V2C-Animation is lower (0.79), which may be due to the influence of noise and background music. In contrast, the GRID dataset is recorded in a studio environment. (2) Changing the measurement metric has relatively little impact on the GRID dataset. V2C has a very obvious decline in WavLM-TDNN similarity, which is not captured by the GE2E model (GE2E score can reach more than 80%). In the future, we will investigate more robust timbre extractors and use the denoising diffusion probabilistic models to further improve the generated wave quality.